OpenStudy (anonymous):

evaluate the indefinite integral ∫(x-2)e^(x^2-4x-3)dx

ganeshie8 (ganeshie8):

Let \(u=x^2-4x-3\)

OpenStudy (anonymous):

i dont understand how to do the u sustitution

OpenStudy (anonymous):

@ganeshie8 is this correct?

OpenStudy (irishboy123):

what about \(u = (x-2)^2\)

ganeshie8 (ganeshie8):

differentiate your answer and see if you get back the integrand

ganeshie8 (ganeshie8):

\[\dfrac{d}{dx} \left(\color{red}{\dfrac{1}{2}e^{x^2-4x-3}+C}\right) = ?\]

OpenStudy (anonymous):

xe^(x^2)-4

OpenStudy (anonymous):

du is 2x-4.... what do i do with the (X-2) then?

OpenStudy (perl):

2x -4 = 2( x -2 )

OpenStudy (anonymous):

2e^(x^2-4x-3)+C

OpenStudy (perl):

u= x^2-4x-3 du = (2x -4) dx du = 2( x-2) dx du /2 = (x-2) dx

OpenStudy (perl):

$$ \large \int (x-2) e^{x^2-4x-3} dx \\ = \large \int e^{x^2-4x-3} (x-2)dx \\ = \large \int e^{u}~ \frac{du}{2} = \frac 1 2 \large \int e^{u}~ du $$

OpenStudy (anonymous):

1/2 e^(x^2-4x-3) + C

OpenStudy (perl):

thats correct